# Distributionally Robust Receive Beamforming

### v2

@@ III-A3 Role of Channel Estimation
Eqs. (7) and (11) imply that channel estimation is not a necessary step in receive beamforming. The only necessary element, from the perspective of statistical machine learning, is the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of the received signal $\mathbf{x}$ and the transmitted signal $\mathbf{s}$. Therefore, the following two points can be highlighted.
  1. 1)
@@
  2. 2)
    If the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is (or assumed to be) Gaussian, we just need to learn the covariance matrices $\bm{R}_{xs}$ and $\bm{R}_{x}$; cf. (13). If, further, the channel matrix $\bm{H}$ is known, $\bm{R}_{xs}$ and $\bm{R}_{x}$ can be expressed using $\bm{H}$.
@@ III-B Distributional Uncertainty and Distributional Robustness
For ease of conceptual illustration, we start with the following stationary-channel assumption in this subsection: The channel statistics remain unchanged within the communication frame so that the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is fixed over time. That is, pilot data $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$ and non-pilot communication data are drawn from the same unknown distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. For the general case where the channel is not statistically stationary within a frame, see Appendix C of the online supplementary materials; the statistical non-stationarity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ may be due to the time-selectivity of the transmit power matrix $\bm{R}_{s}$, of the channel matrix $\bm{H}$, and/or of the channel noise covariance $\bm{R}_{v}$.
@@ III-B1 Issue of Distributional Uncertainty
denote the empirical distribution supported on the $L$ collected data $\{(\bm{x}_{i},\bm{s}_{i})\}_{i\in[L]}$, where $\delta_{(\bm{x}_{i},\bm{s}_{i})}$ denotes the Dirac distribution (i.e., point-mass distribution) centered on $(\bm{x}_{i},\bm{s}_{i})$; note that ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is a discrete distribution. If we use the estimated joint distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ as a surrogate of the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, (7) becomes the conventional empirical risk minimization (ERM) $\min_{\bm{\phi}\in\mathcal{B}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$, i.e.,
@@
There exists the distributional difference between the sample-defined nominal distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and true data-generating distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ due to the limited size of the training data set (i.e., limited pilot length) and the time-selectivity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. From the perspective of applied statistics and machine learning, the distributional difference between ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ (i.e., the distributional uncertainty of ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ compared to $\mathbb{P}_{\mathbf{x},\mathbf{s}}$) may cause significant performance degradation of (15) compared to (7), so is the performance deterioration of (16) compared to (12). For extensive reading on this point, see Appendix C of the online supplementary materials. Therefore, to reduce the adverse effect introduced by the distributional uncertainty in ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$, a new surrogate of (7) rather than the sample-averaged approximation in (15) is expected.
III-B2 Distibutionally Robust Estimation
@@
$\mathcal{U}_{\mathbf{x},\mathbf{s}}\coloneqq\{\mathbb{P}_{\mathbf{x},\mathbf{s}}|d(\mathbb{P}_{\mathbf{x},\mathbf{s}},\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}})\leq\epsilon\},$ (18)
@@
If $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ contains only Gaussian distributions, (17) is particularized to
@@ IV-A General Framework and Concrete Examples
Suppose that the constraint $d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0}$ is compact convex and $\bm{R}_{x}$ is invertible. Let $\bm{R}^{\star}$ solve the problem below:
@@
Then $(\bm{W}^{\star},\bm{R}^{\star})$ is a solution to Problem (19). On the other hand, if $(\bm{W}^{\prime},\bm{R}^{\prime})$ solves Problem (19), then $\bm{R}^{\prime}$ is a solution to (21) and $(\bm{W}^{\prime},\bm{R}^{\prime})$ satisfies (22).
@@
See Appendix D of the online supplementary materials. $\square$
@@
See Appendix E of the online supplementary materials. $\square$
@@
$\hat{\bm{R}}-\epsilon_{0}\bm{E}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{E},$ (25)
for some $\bm{E}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$ such that $\hat{\bm{R}}-\epsilon_{0}\bm{E}\succeq\bm{0}$. $\square$
@@
$\hat{\bm{R}}-\epsilon_{0}\bm{I}_{N+M}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{I}_{N+M},$ (26)
for some $\epsilon_{0}\geq 0$ such that $\hat{\bm{R}}-\epsilon_{0}\bm{I}_{N+M}\succeq\bm{0}$. $\square$
@@
The generalized diagonal-loading uncertainty set of $\bm{R}$ is constructed as
@@
for some $\bm{F},\bm{G}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$ such that the matrix in the first line is positive semi-definite. $\square$
@@
$\theta_{1}\hat{\bm{R}}\preceq\bm{R}\preceq\theta_{2}\hat{\bm{R}},$ (28)
@@ 1.  C1)
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-AM}}&=(\hat{\bm{R}}_{xs}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{E}_{x})^{-1}\\ &=(\hat{\bm{H}}\hat{\bm{R}}_{s}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}\cdot\\ &\quad\quad\quad[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{0}\bm{E}_{x}]^{-1}.\end{array}$ (29)
@@ 2.  C2)
    Under Definition 2, the diagnal-loading distributionally robust (DR-DL) beamformer is
@@ 4.  C4)
$\hat{\bm{R}}_{x}-\epsilon_{0}\bm{F}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{F}.$ (33)
@@
$\hat{\bm{R}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{I}_{N},$ (34)
@@
As we can see from Corollary 1, the primary benefit of using the moment-based uncertainty sets is the computational simplicity due to the availability of closed-form solutions. If the uncertainty sets are constructed using the Wasserstein distance $\sqrt{\operatorname{Tr}[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}]}\leq\epsilon_{0}$ or the KL divergence $\frac{1}{2}[\operatorname{Tr}[{\hat{\bm{R}}^{-1}\bm{R}-\bm{I}_{N+M}}]-\ln\det(\hat{\bm{R}}^{-1}\bm{R})]\leq\epsilon_{0}$ between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$, the induced distributionally robust linear beamforming problems have no closed-form solutions, and therefore, are computationally prohibitive in practice. In addition, Corollary 1 suggests that the distributionally robust beamformer under the multiplicative moment uncertainty set (28) is the same as the nominal beamformer $\hat{\bm{R}}^{\mathsf{H}}_{xs}\hat{\bm{R}}_{x}^{-1}$, which essentially do not introduce robustness in wireless signal estimation; this is another motivation why we construct new moment-based uncertainty sets in Definitions 1, 2, and 3. However, we can modify the multiplicative moment uncertainty set in Definition 4 to achieve robustness.
@@
The modified multiplicative moment uncertainty set of $\bm{R}$ is given as
@@
The robust beamformer under the modified multiplicative moment uncertainty set (35) is $\bm{W}^{\star}_{\text{DR-MMM}}=\hat{\bm{R}}^{\mathsf{H}}_{xs}\cdot[\theta_{2}\hat{\bm{R}}_{x}]^{-1}$.
@@
which is equivalent, in the sense of the same minimizer, to
@@
$\hat{\bm{R}}_{x,\text{thr}}\coloneqq\bm{Q}{\footnotesize\left[\begin{array}[]{cccc}\lambda_{1}&&&\\ &\max\{\mu\lambda_{1},\lambda_{2}\}&&\\ &&\ddots&\\ &&&\max\{\mu\lambda_{1},\lambda_{N}\}\end{array}\right]}\bm{Q}^{-1},$ (39)
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR}}=[\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\epsilon_{1}\bm{H}^{\mathsf{H}}(\bm{H}\bm{H}^{\mathsf{H}})^{-1}]\cdot\\ \quad\quad\quad\quad\quad[\bm{H}\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\epsilon_{1}\bm{I}_{M}]^{-1},\end{array}$ (43)
@@
Consider a linear regression problem
@@
and the distributionally robust linear estimator of $\bm{W}$
@@
then the distributionally robust linear estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method
@@
and the distributionally robust linear estimator of $\bm{W}$
@@
Suppose that $\mathbf{\Delta}_{x}$ is uncorrelated with $\mathbf{x}$, with $\mathbf{s}$, and with $\mathbf{\Delta}_{s}$; in addition, $\mathbf{\Delta}_{s}$ is uncorrelated with $\mathbf{x}$. If the second-order moment of $\mathbf{\Delta}_{x}$ is upper bounded as $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon_{0}\bm{I}_{N}$, then the distributionally robust linear estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method

### v3

@@ III-A3 Role of Channel Estimation
Eqs. (7) and (10) imply that channel estimation is not a necessary step in receive combining. The only necessary element, from the perspective of statistical machine learning, is the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of the received signal $\mathbf{x}$ and the transmitted signal $\mathbf{s}$. Therefore, the following two points can be highlighted.
  1. a)
@@
  2. b)
    If the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is (or assumed to be) Gaussian, we just learn covariance matrices $\bm{R}_{xs}$ and $\bm{R}_{x}$; cf. (12); Gaussianity assumption of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is beneficial in reducing computational burdens. If, further, the channel matrix $\bm{H}$ is known, $\bm{R}_{xs}$ and $\bm{R}_{x}$ can be expressed using $\bm{H}$.
@@ III-B Distributional Uncertainty and Distributional Robustness
For ease of conceptual illustration, we start with the following stationary-channel assumption in this subsection: The channel statistics remain unchanged within the communication frame so that the joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is fixed over time. That is, pilot data $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$ and non-pilot communication data are drawn from the same unknown distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. For the general case where the channel is not statistically stationary within a frame, see Appendix C; the statistical non-stationarity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ may be due to the time-selectivity of the transmit power matrix $\bm{R}_{s}$, of the channel matrix $\bm{H}$, and/or of the channel noise covariance $\bm{R}_{v}$.
@@ III-B1 Issue of Distributional Uncertainty
denote the empirical distribution supported on the $L$ collected data $\{(\bm{x}_{i},\bm{s}_{i})\}_{i\in[L]}$, where $\delta_{(\bm{x}_{i},\bm{s}_{i})}$ denotes the Dirac distribution (i.e., point-mass distribution) centered on $(\bm{x}_{i},\bm{s}_{i})$; note that ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is a discrete distribution. If we use the estimated joint distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ as a surrogate of the true joint distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$, (7) becomes the conventional empirical risk minimization (ERM)
$\min_{\bm{\phi}\in\mathcal{B}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}},$ (14)
i.e.,
@@
There exists the distributional difference between the sample-defined nominal distribution ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and true data-generating distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ due to the limited size of the training data set (i.e., limited pilot length) and the time-selectivity of $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. From the perspective of applied statistics and machine learning, the distributional difference between ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ (i.e., the distributional uncertainty of ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ compared to $\mathbb{P}_{\mathbf{x},\mathbf{s}}$) may cause significant performance degradation of (15) compared to (7), so is the performance deterioration of (16) compared to (11). For extensive reading on this point, see Appendix C. Therefore, to reduce the adverse effect introduced by the distributional uncertainty in ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$, a new surrogate of (7) rather than the sample-averaged approximation in (15) is expected.
III-B2 Distributionally Robust Estimation
@@
$\mathcal{U}_{\mathbf{x},\mathbf{s}}\coloneqq\{\mathbb{P}_{\mathbf{x},\mathbf{s}}|~{}d(\mathbb{P}_{\mathbf{x},\mathbf{s}},\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}})\leq\epsilon\},$ (18)
@@
If $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ contains (or is assumed, for computational simplicity, to contain) only Gaussian distributions, (17) particularizes to
@@
For additional discussions on the framework of distributionally robust estimation, see Appendix D.
@@ IV-A General Framework and Concrete Examples
Suppose that the set $\{\bm{R}|~{}d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0}\}$ is compact convex and $\bm{R}_{x}$ is invertible. Let $\bm{R}^{\star}$ solve the problem below:
@@
Then $(\bm{W}^{\star},\bm{R}^{\star})$ is a solution to Problem (19). On the other hand, if $(\bm{W}^{\star},\bm{R}^{\star})$ solves Problem (19), then $\bm{R}^{\star}$ is a solution to (21) and $(\bm{W}^{\star},\bm{R}^{\star})$ satisfies (22).
@@
See Appendix E. $\square$
@@
See Appendix F. $\square$
@@
$\{\bm{R}|~{}\hat{\bm{R}}-\epsilon_{0}\bm{E}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{E},~{}\bm{R}\succeq\bm{0}\}$ (25)
for some $\bm{E}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$. $\square$
@@
$\{\bm{R}|~{}\hat{\bm{R}}-\epsilon_{0}\bm{I}_{N+M}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{I}_{N+M},~{}\bm{R}\succeq\bm{0}\}$ (26)
for some $\epsilon_{0}\geq 0$. $\square$
@@
The generalized diagonal-loading uncertainty set of $\bm{R}$ is constructed by the following constraints: $\bm{R}\succeq\bm{0}$ and
@@
for some $\bm{F},\bm{G}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$. $\square$
@@
$\{\bm{R}|~{}\theta_{1}\hat{\bm{R}}\preceq\bm{R}\preceq\theta_{2}\hat{\bm{R}}\}$ (28)
@@ 1.  C1)
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-AM}}&=(\hat{\bm{R}}_{xs}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{E}_{x})^{-1}\\ &=(\hat{\bm{H}}\hat{\bm{R}}_{s}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}\cdot\\ &\quad\quad\quad[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{0}\bm{E}_{x}]^{-1},\end{array}$ (29)
    where $\hat{\bm{H}}$, $\hat{\bm{R}}_{s}$, and $\hat{\bm{R}}_{v}$ denote the estimates of $\bm{H}$, $\bm{R}_{s}$, and $\bm{R}_{v}$, respectively.
@@ 2.  C2)
    Under Definition 2, the diagonal-loading distributionally robust (DR-DL) beamformer is
@@ 4.  C4)
$\{\bm{R}_{x}|~{}\hat{\bm{R}}_{x}-\epsilon_{0}\bm{F}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{F},~{}\bm{R}_{x}\succeq\bm{0}\}.$ (33)
@@
$\{\bm{R}_{x}|~{}\hat{\bm{R}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{I}_{N},~{}\bm{R}_{x}\succeq\bm{0}\},$ (34)
@@
As we can see from Corollary 1, the primary benefit of using the moment-based uncertainty sets is the computational simplicity due to the availability of closed-form solutions. If the uncertainty sets are constructed using the Wasserstein distance $\sqrt{\operatorname{Tr}[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}]}\leq\epsilon_{0}$ or the KL divergence $\frac{1}{2}[\operatorname{Tr}[{\hat{\bm{R}}^{-1}\bm{R}-\bm{I}_{N+M}}]-\ln\det(\hat{\bm{R}}^{-1}\bm{R})]\leq\epsilon_{0}$ between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$, the induced distributionally robust linear beamforming problems have no closed-form solutions, and therefore, are computationally prohibitive in practice. In addition, Corollary 1 suggests that the distributionally robust beamformer under the multiplicative moment uncertainty set (28) is the same as the nominal beamformer $\hat{\bm{R}}^{\mathsf{H}}_{xs}\hat{\bm{R}}_{x}^{-1}$, which essentially does not introduce robustness in wireless signal estimation; this is another motivation why we construct new moment-based uncertainty sets in Definitions 1, 2, and 3. However, we can modify the multiplicative moment uncertainty set in Definition 4 to achieve robustness.
@@
The modified multiplicative moment uncertainty set of $\bm{R}$ is defined by the following constraint:
@@
The robust beamformer under the modified multiplicative moment uncertainty set (35) is
$\bm{W}^{\star}_{\text{DR-MMM}}=\hat{\bm{R}}^{\mathsf{H}}_{xs}\cdot[\theta_{2}\hat{\bm{R}}_{x}]^{-1}.$ (36)
@@
which is equivalent, in the sense of the same solutions, to
@@
$\begin{array}[]{l}\hat{\bm{R}}_{x,\text{thr}}\coloneqq\\ \quad\bm{Q}\left[\begin{array}[]{cccc}\lambda_{1}&&&\\ &\max\{\mu\lambda_{1},\lambda_{2}\}&&\\ &&\ddots&\\ &&&\max\{\mu\lambda_{1},\lambda_{N}\}\end{array}\right]\bm{Q}^{-1},\end{array}$ (40)
@@
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR}}=[\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\epsilon_{1}\bm{H}^{\mathsf{H}}(\bm{H}\bm{H}^{\mathsf{H}})^{-1}]\cdot\\ \quad\quad\quad\quad\quad[\bm{H}\hat{\bm{R}}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\epsilon_{1}\bm{I}_{N}]^{-1},\end{array}$ (44)
@@
Consider a linear regression problem on $(\mathbf{x},\mathbf{s})$, i.e.,
@@
where $\mathbf{e}$ denotes the error term and the distributionally robust estimator of $\bm{W}$, i.e.,
@@
then the distributionally robust estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method
@@
and the distributionally robust estimator of $\bm{W}$
@@
Suppose that $\mathbf{\Delta}_{x}$ is uncorrelated with $\mathbf{x}$, with $\mathbf{s}$, and with $\mathbf{\Delta}_{s}$; in addition, $\mathbf{\Delta}_{s}$ is uncorrelated with $\mathbf{x}$. If the second-order moment of $\mathbf{\Delta}_{x}$ is upper bounded as $\mathbb{E}\mathbf{\Delta}_{x}\mathbf{\Delta}^{\mathsf{H}}_{x}\preceq\epsilon_{0}\bm{I}_{N}$, then the distributionally robust estimator of $\bm{W}$ becomes a ridge regression (i.e., squared-$F$-norm regularized) method